# On $[1,2]$-Domination in Interval and Circle Graphs

### v2

@@ 3 $[1,2]$-domination in Circle Graphs
Proof.
Let $D$ of size $k=7m+2t$ be the given $[1,2]$-dominating set. We show there is a truth assignment to the variables in $X$ that satisfy $I$. Recall that associated with each literal $l$ that appears in a clause $C_{j}$, there is a pair of chords $p_{j}^{l}$. Since no chord is adjacent to two different pairs $p$’s, $D$ must contain, for each $l$ and $j$, either $t_{j}^{l}$ or $f_{j}^{l}$ or both chords in $p$. If for each literal $l$ in a clause $j$, $D$ contains $t_{j}^{l}$ or $f_{j}^{l}$(but not both) then exactly $3m$ type $t$ or $f$ chord in $D$ suffice to dominate all pairs of chords $p$. Otherwise, if $D$ contains any pairs of $p$ more than $3m$ are necessary to dominate chords in pairs $p$. Also, all $c_{1}$’s and $c_{2}$’s, totally $2m$, chords are in $D$ to dominate the three crossed chord to them.
There is no chord in $G(I)$ adjacent to $(u,w)$ pairs associated with different clauses. Thus the set of chords in $D$ which dominate the $(u,w)$ pairs in different clauses are pairwise disjoint. Also, no chord in $G(I)$ is adjacent to all the three pairs $(u,w)$ in a clause. Therefore $D$ must contain at least two chords per a clause to dominate all $(u,w)$ pairs in a clause. Combined with the argument in the previous paragraph, this means exactly $3m$ chords in $D$ dominate $p$ pairs and $2m$ chords dominate the $(u,w)$ pairs. This further implies that for each variable $x_{i}$ appearing in a clause $C_{j}$, either $t_{j}^{l}$ or $f_{j}^{l}$ belongs to $D$.
Suppose three literal $l_{1},l_{2}$, and $l_{3}$ in clause $C_{j}$, and assume none of $t_{j}^{1},t_{j}^{2}$, and $t_{j}^{3}$ contain in $D$ (All $f$ chords are selected). So, all pairs $a_{j}^{1},a_{j}^{2}$ and $a_{j}^{3}$ are not dominated. There is no $(u,w)$ pair that can dominate all three pairs $a_{j}$’s. So at least one $t$ chord must be selected to dominate one $a_{j}$ pair, and the other remaining $a_{j}$’s will be dominated by the appropriate pair $(u,w)$.
Till now, we show that for each clause $C_{j}$, $D$ contains at least one $t$ chord and one pair $(u,w)$. Based on the construction of the gadget, each chord $tt,tf,ft$, and $ff$ are attached to a claw chord, centered at $s$. So, all the chords $s$ must be in domination and as we have $2t$ such chords, then the total number of our $[1,2]$-domination is equal to $|D|=7m+2t$. Now, we can construct a satisfying truth assignment for $I$ as follows. If $D$ contains a $t$ chord associated with a literal involving $x_{i}$, then we set this literal(and the corresponding value to its variable) to true in $I$. Similarly, $D$ contains a $f$ chord associated with a literal involving $x_{i}$, then we set this literal to false in $I$. ∎
The proof of theorem 2 concludes from Lemma 10 and 11. ∎
@@ 4 Conclusion
Applying a constraint that determines the maximum number of times a vertex outside the dominating set, like $D$, is dominated by the elements inside $D$, makes the problem easier or harder than the original domination problem. In this paper, first we have designed a polynomial algorithm for $[1,2]$-dominating problem in proper interval graphs. Although it seems $O(n^{4})$ is huge, it is an affirmative answer to the open problem on the complexity class of $[1,2]$-Domination problem in non-proper interval graphs. Designing more efficient algorithm for this problem is an interesting open problem. Then, we have shown that the $[1,2]$-Dominating Set problem is $NP$-complete on circle graphs. It is thus natural to ask for which subclass of circle graphs, like distance-hereditary or k-polygon, the problem has polynomial-time algorithm. Meanwhile, it would be desirable to show the complexity of $[1,2]$-dominating problem on other classes of graphs marked with a question mark in Figure 3.

### v4

@@ 3 $[1,2]$-domination in Circle Graphs
Proof 3.21.
Let $D$ of size $k=7m+2t$ be the given $[1,2]$-dominating set. We show there is a truth assignment to the variables in $X$ that satisfies $I$. Recall that associated with each literal $l$ that appears in a clause $C_{j}$, there is a pair of chords $p_{j}^{l}$. Since no chord is adjacent to two different pairs $p$’s, $D$ must contain, for each $l$ and $j$, either $t_{j}^{l}$ or $f_{j}^{l}$ or both chords in $p$. If for each literal $l$ in a clause $j$, $D$ contains $t_{j}^{l}$ or $f_{j}^{l}$(but not both), then exactly $3m$ type $t$ or $f$ chord in $D$ are sufficient to dominate all pairs of chords $p$. Otherwise, if $D$ contains any additional pairs of $p$ more than $3m$ chords are necessary to dominate chords in pairs $p$. Also, all $c_{1}$’s and $c_{2}$’s, totally $2m$, chords are in $D$ to dominate the three crossed chords connected to them.
There is no chord in $G(I)$ adjacent to $(u,w)$ pairs associated with different clauses. Thus the set of chords in $D$ which dominate the $(u,w)$ pairs in different clauses are pairwise disjoint. Also, no chord in $G(I)$ is adjacent to all the three pairs $(u,w)$ in a clause. Therefore $D$ must contain at least two chords per clause to dominate all $(u,w)$ pairs in a clause. Combined with the argument in the previous paragraph, this means exactly $3m$ chords in $D$ dominate $p$ pairs and $2m$ chords dominate the $(u,w)$ pairs. This further implies that for each variable $x_{i}$ appearing in a clause $C_{j}$, either $t_{j}^{l}$ or $f_{j}^{l}$ belongs to $D$.
Suppose there are three literal $l_{1},l_{2}$, and $l_{3}$ in clause $C_{j}$, and assume none of $t_{j}^{1},t_{j}^{2}$, and $t_{j}^{3}$ are included in $D$ (i.e., all $f$ chords are selected). In this case, all pairs $a_{j}^{1},a_{j}^{2}$, and $a_{j}^{3}$ are not dominated. There is no single $(u,w)$ pair that can dominate all three pairs $a_{j}$’s. So at least one $t$ chord must be selected to dominate one $a_{j}$ pair, and the other remaining $a_{j}$’s will be dominated by the appropriate pair $(u,w)$.
Until now, we have shown that for each clause $C_{j}$, $D$ contains at least one $t$ chord and one pair $(u,w)$. Based on the construction of the gadget, each chord $tt,tf,ft$, and $ff$ are attached to a claw chord, centered at $s$. So, all the chords $s$ must be in domination and as we have $2t$ such chords, then the total size of our $[1,2]$-domination is equal to $|D|=7m+2t$. Now, we can construct a satisfying truth assignment for $I$ as follows. If $D$ contains a $t$ chord associated with a literal involving $x_{i}$, then we set this literal(and the corresponding value to its variable) to true in $I$. Similarly, $D$ contains a $f$ chord associated with a literal involving $x_{i}$, then we set this literal to false in $I$.
The proof of theorem 3.16 follows directly from Lemma 3.18 and 3.20.
@@ 4 Conclusion
Applying a constraint that determines the maximum number of times a vertex outside the dominating set, like $D$, is dominated by the elements inside $D$, can either make the problem easier or harder than the domination problem. In this paper, we first designed a polynomial algorithm for the $[1,2]$-dominating problem in proper interval graphs. Although $O(n^{4})$ might seem large, it provides an affirmative answer to the open problem regarding the complexity class of the $[1,2]$-Domination problem in non-proper interval graphs. Designing a more efficient algorithm for this problem remains an interesting open problem. Next, we have shown that the $[1,2]$-Dominating Set problem is $NP$-complete on circle graphs. It is thus natural to ask for which subclass of circle graphs, like distance-hereditary or k-polygon, the problem has a polynomial-time algorithm. Meanwhile, it would be desirable to show the complexity of $[1,j]$-dominating problem on other classes of graphs marked with a question mark in Figure 3.